# Elastic Metastability: Thermally Activated Snap-through Transitions in Nanostructures

### v1

@@ Abstract
The morphological snap-through transition of a bistable elastic structure typically requires external loading, unless it is miniaturized to the nanoscale, where thermal fluctuations play a significant role. We propose that metastability can emerge in microscopic elastic systems when coupled to a thermal environment, giving rise to thermally activated transitions between metastable configurations. Using atomistic simulations combined with rare event methods, we demonstrate that snap-through events can occur as well-defined thermally activated transitions in a geometrically constrained graphene nanoribbon. Well-tempered metadynamics is employed to determine the transition energetics and pathways. Notably, the temperature dependence of the transition rate constant is accurately described by generalized transition state theory in terms of the Landau free energy. This study introduces a theoretical framework for understanding elastic metastability, extends reaction rate theory to nanomechanical systems, and suggests strategies for designing temperature-responsive nanodevices.
preprint: APS/123-QED
@@
Thermally activated transitions underlie a great variety of processes in physics, chemistry, and biology, with examples ranging from magnetization reversal [1, 2] and solid state diffusion [3] to chemical reactions [4, 5] and protein folding [6]. A common characteristic of these systems is the presence of a free energy landscape with metastable states, between which transitions are driven by thermal fluctuations. In the context of elastic systems, although thermal fluctuations are not accounted for in a macroscopic elastic continuum, they may play an essential role in the microscopic scenario. Yet to date, thermally activated morphological transitions have neither been observed nor predicted in microscopic elastic systems. Therefore, we propose that, when miniaturized to the nanoscale, certain bistable elastic structures can undergo morphological transitions via thermal activation.
@@
A bistable elastic structure possesses two stable equilibrium states. A morphological snap-through transition occurs when such a structure rapidly switches from one equilibrium state to the other. These transitions allow for the generation of fast movements through the release of stored elastic energy and are manifested in various biological and engineered systems, including the Venus flytrap [7], hummingbird beaks [8], mechanical metamaterials [9], and nanoelectromechanical switches [10]. In macroscopic systems, a snap-through transition is typically realized either by applying an external load that overcomes the energy barrier, or by imposing boundary actuation that destabilizes one of the equilibrium states [11]. However, due to the presence of thermal fluctuations in microscopic systems, such external interventions may not be required for a snap-through transition to occur. If a nanoscale bistable structure exhibits an energy barrier that can be surmounted by thermal fluctuations on physically relevant timescales, snap-through transitions would occur spontaneously with an associated rate constant, rendering the system metastable. While previous computational studies have investigated the snap-through of buckled graphene membranes [12, 13], the role of thermal activation in driving such transitions at the nanoscale remains unexplored, because the systems studied were subject to prohibitively large energy barriers precluding any spontaneous transition. In this work, we consider a buckled graphene nanoribbon (GNR) under geometric confinement, in which the energy barrier can be modulated to permit thermal activation.
@@
The GNR can serve as a nanoscale counterpart of elastic strips, which have been increasingly adopted as a canonical system for analyzing the dynamics and mechanisms of snap-through transitions in bistable structures [14, 15, 16, 17, 18, 19, 20, 21]. Here, we consider a zigzag GNR with a length of 15.23 nm and N = 16 zigzag chains, whose edges are hydrogen-terminated to eliminate intrinsic mechanical instabilities [22]. When the GNR is placed under appropriate geometric constraints, in the same manner as a buckled elastic strip with clamped boundary conditions [19], it buckles into one of two stable configurations. Given the large length and time scales involved in the processes of interest, molecular dynamics (MD) simulations with empirical interatomic potentials constitute the only viable and effective computational approach that retains atomistic resolution. Empirical interatomic potentials have been widely used to describe the mechanical and thermodynamic properties of two-dimensional materials [23]. Among them, the second-generation reactive empirical bond-order (REBO-II) potential [24] is one of the most commonly used for solid carbon and hydrocarbon systems, and has been shown to reliably capture the elasticity [25] and thermomechanical behaviors [26] of graphene. Accordingly, we perform MD simulations for the H-terminated GNR system described by the REBO-II potential to investigate its thermally activated snap-through transitions. Furthermore, rare event methods, including enhanced sampling methods and reaction rate theory, are employed to unravel the free energy landscape and transition kinetics.
Figure 1: (a) Representative configurations of the GNR corresponding to the natural state (top), the asymmetric unstable equilibrium ($\mathcal{A}$ state, middle), and the inverted state (bottom). The rigidly constrained atoms at the two ends impose a fixed end-shortening $\Delta L$ and a symmetric tangent angle $\alpha$ at the boundaries. (b) The bifurcation diagram obtained from the Euler–beam model, with solid lines indicating the natural state (violet upper branch) and the inverted state (blue lower branch), and the orange dashed line indicating the $\mathcal{A}$ state. Diamond markers represent the equilibrium averages of $w_{0}$ from MD simulations.
@@
The GNR can be transformed into a bistable structure when subject to geometric constraint at the boundaries. To achieve this, the outermost seven rows of atoms at each end are rigidly held, leaving a flexible segment of length $L$ = 13.75 nm. A fixed end-shortening $\Delta L=0.08L$ is applied between the two ends, which are then rotated in opposite directions by a small adjustable angle $\alpha$, causing the GNR to adopt one of two mechanical stable states: a natural state and an inverted state, as shown in Fig. 1(a). The equilibrium configurations of these states can be analyzed with Euler–Bernoulli beam theory [14, 19, 21] in the limit of small transverse displacement $w(x,t)$,
@@
where $\rho$, $b$, and $h$ are the effective density, width, and thickness of the GNR, respectively; $B$ denotes the bending stiffness, and $F$ the applied compressive force. The geometric constraint enters the model through boundary conditions and the end-shortening confinement condition [14]. The time-independent solutions of Eq. (1) capture the equilibrium behavior of the system, which can be summarized in a bifurcation diagram in Fig. 1(b). The midpoint displacement $w_{0}$ of the GNR, weighted by $L$, is plotted as a function of the dimensionless control parameter $\mu=\alpha\sqrt{L/\Delta L}$. As $\mu$ increases, the inverted state loses its stability at $\mu^{*}=2$, where it intersects an asymmetric unstable equilibrium branch (denoted as $\mathcal{A}$ state) in a subcritical pitchfork bifurcation. In a macroscopic system, $\mu^{*}$ corresponds to the onset of snap-through towards the natural state. As the inverted state further approaches $\mu^{*}_{2}\approx 2.012$, it undergoes a saddle-node bifurcation and vanishes. Meanwhile, the natural state persists stably over the full range of $\mu$.
To compare the behavior of the GNR in atomistic simulations with the description of the Euler-beam equation, we monitor $w_{0}$ of the GNR at different values of $\mu$, by setting $\alpha$ to successive integer degrees. For each $\alpha$, the GNR is fully thermalized in the natural and the inverted states separately, and a 5 ns trajectory is initiated from each state to calculate the time average of $w_{0}$. A Langevin thermostat is used to provide the thermalizing environment at 300 K. The time-averaged $w_{0}$ from our MD simulations is found to be in good agreement with the Euler–beam model, as shown in Fig. 1(b). Naturally, the GNR configuration does not remain stationary at the mechanical equilibrium position of the natural or inverted state, but instead fluctuates around it, thereby defining a thermodynamic equilibrium state. When $\alpha$ exceeds $30^{\circ}$ ($\mu\approx 1.851$), the GNR cannot persist in the vicinity of the inverted state throughout a 5 ns trajectory—it snaps to the natural state under the influences of thermal fluctuations that are not captured by Eq. (1). This indicates that the thermal stability of the inverted state has already been compromised, even though the system is well below $\mu^{*}$, the threshold of mechanical instability.
@@
where $\delta w_{0}(t)=w_{0}(t)-\langle w_{0}\rangle$, and $\langle\cdot\rangle$ denotes the ensemble average in the inverted state. The profile of $\widetilde{C}(t)$ is presented in Fig. 2(b). The time integral of $\widetilde{C}(t)$ yields the dominant local relaxation time scale $\tau_{s}$ = 0.32 ns. In addition, we analyze the distribution of escape times in the 100 trajectories. The mean escape time is estimated using maximum likelihood estimation (MLE), yielding a value of $\tau_{e}=51.70$ ns, which is two orders of magnitude greater than the local relaxation time $\tau_{s}$. The cumulative distribution function (CDF) of the escape times is shown in Fig. 2(c), along with the fitted exponential distribution obtained from MLE. A Kolmogorov–Smirnov test produces a $p$-value of 0.957, confirming that the escape times follow an exponential distribution. The separation of time scales ($\tau_{e}\gg\tau_{s}$) [4, 5, 27] and the exponential distribution of escape times each demonstrate that the snap-through transitions in the GNR constitute a well-defined thermally activated rate process.
@@
In Fig. 2(a), even while remaining within the local equilibrium states, the system exhibits substantial fluctuations in the atomistic potential energy, which obscures a clear description of the transition energetics. Therefore, instead of relying on elastic energy, a free-energy-based approach becomes necessary to capture the thermally activated snap-through at the nanoscale. Here, to determine the free energy profile of the system, we employ well-tempered metadynamics (WT-MetaD) [28], an enhanced sampling method that has been extensively used in complex molecular systems to accelerate rare events and accurately calculate free energy [29, 30].
Figure 3: (a) Free energy profile computed from WT-MetaD (five 500 ns production runs) along the order parameter $w_{0}$ for the GNR constrained at $\alpha=29^{\circ}$ ($\mu\approx 1.789$). Inset: Magnified view of the free energy profile highlighting the forward transition barrier $F^{+}$. (b) Forward and backward transition barriers, $F^{+}$ and $F^{-}$, as functions of the dimensionless geometric parameter $\mu$.
Because $w_{0}$ is the natural order parameter of the snap-through transition, one can perform one-dimensional WT-MetaD simulations using $w_{0}$ as the biased collective variable. At $\alpha=29^{\circ}$, the free energy profile $F(w_{0})$ computed from WT-MetaD is presented in Fig. 3(a), exhibiting two local minima corresponding to the inverted state and the natural state. The forward transition from the inverted to the natural state is associated with a free energy barrier of $F^{+}$ = 71.823 meV, while the backward transition corresponds to a much higher barrier of $F^{-}$ = 2526.776 meV. The substantial height of $F^{-}$ suggests that the reverse transition is kinetically suppressed, as will be addressed in our later analysis. Since the energetics of the system can be controlled by the geometric parameter $\mu$, we have further computed the free energy barriers from WT-MetaD at various values of $\mu$, as shown in Fig. 3(b). When $\mu=0$, the forward and backward transition barriers are equal, implying no energetic preference between the two states. As $\mu$ increases, the barriers become increasingly unequal.
Figure 4: Free energy surface computed from two-dimensional WT-MetaD in the ($w_{0}$, $\Delta w_{\text{qrt}}$) space. The collective variable $\Delta w_{\text{qrt}}$ quantifies the degree of asymmetry in the GNR. The locations of the natural state (red circle), inverted state (blue diamond), and the $\mathcal{A}$ state (yellow triangle), obtained from the Euler-beam model, coincide with the metastable and transition states identified on the FES. For illustrative clarity, the FES is shown for $\alpha=16^{\circ}$ ($\mu\approx 0.9873$).
@@
To explore the free energy landscape of the system, we extend our approach to two-dimensional WT-MetaD involving two collective variables. Motivated by the important role of symmetry breaking in bistable elastic strips [19, 21], we propose a measure of asymmetry that is readily amenable to biased sampling in WT-MetaD, defined as the difference in transverse displacement between the quarter-length positions measured from the two ends of the GNR: $\Delta w_{\text{qrt}}=w(L/4)-w(-L/4)$. The WT-MetaD simulations are then performed by biasing both $w_{0}$ and $\Delta w_{\text{qrt}}$, resulting in the free energy surface (FES) in the ($w_{0}$, $\Delta w_{\text{qrt}}$) space, as shown in Fig. 4. Two basins corresponding to the inverted state and the natural state are identified, with local minima aligned with $\Delta w_{\text{qrt}}=0$, indicative of their symmetric nature. The transition pathways between the two states are characterized by the minimum free energy paths (MFEPs) connecting the basins, along which the system transitions between the two symmetric states via highly asymmetric intermediate states. To compare with the elastic continuum theory, the mechanical equilibrium positions from the Euler-beam model are plotted on the FES, showing good agreement with the metastable minima. Moreover, the asymmetric unstable equilibria ($\mathcal{A}$ state) coincide with a pair of first-order saddle points, indicating their role as the transition state in thermally activated transitions. It can be readily confirmed that the trajectories of one-dimensional WT-MetaD faithfully track the MFEPs on the two-dimensional landscape (see the Supplemental Material). Thus, the resulting free energy profile $F(w_{0})$ unambiguously corresponds to the Landau free energy governing the transitions.
Reaction rate theory has been widely employed to describe the kinetics of activated rate processes [4, 5], prominently including Kramers’ theory [31] and transition state theory (TST) [32]. Here we adopt generalized transition state theory in terms of the potential of mean force [33] to characterize the kinetics of the thermally activated snap-through transition, with emphasis on elucidating the temperature dependence of its transition rate constant.
@@
Provided that there exists a collective variable $\xi$ such that the transition state dividing surface $\xi(\mathbf{r})=\xi^{\ddagger}$ separates the reactant region $\{\mathbf{r}|\xi(\mathbf{r})<\xi^{\ddagger}\}$ from the product region $\{\mathbf{r}|\xi(\mathbf{r})>\xi^{\ddagger}\}$ in configuration space, the transition rate constant can be calculated as the one-directional reactive flux through this surface. For a system characterized by the one-dimensional potential of mean force $W(\xi)$, the generalized TST rate constant takes the form
$k_{\mathrm{TST}}$ $=\frac{\langle{Z_{\xi}}^{\frac{1}{2}}\rangle_{\xi=\xi^{\ddagger}}}{(2\pi\beta)^{\frac{1}{2}}}\frac{\mathrm{e}^{-\beta W(\xi^{\ddagger})}}{\int_{\mathrm{R}}\mathrm{d}\xi\,\mathrm{e}^{-\beta W(\xi)}}$ (3)
@@
where $\langle\cdot\rangle_{\xi=\xi^{\ddagger}}$ denotes the ensemble average sample at the transition state, $\int_{\mathrm{R}}$ denotes the integral over the reactant region, and $\Delta\xi$ is given by $\Delta\xi=\int_{\mathrm{R}}\mathrm{d}\xi$; $\beta$ is the inverse temperature, $\Delta F^{\ddagger}$ is the free energy of activation, and $Z_{\xi}$ is the inverse effective mass associated with $\xi$, defined as
@@
with $M_{i}$ being the atomic mass associated with the $i$th component of $\mathbf{r}$. For the GNR system under investigation, the order parameter $w_{0}$ serves the role of $\xi$, so that the Landau free energy $F(w_{0})$ can be used in place of $W(\xi)$ in Eq. (3).
We justify the applicability of generalized TST to the activated snap-through transition by examining the consistency between the free energy of activation $\Delta F^{\ddagger}$ inferred from Eq. (4) based on the transition rates determined in unbiased simulations at different temperatures and the $\Delta F^{\ddagger}$ calculated directly from $F(w_{0})$ given by WT-MetaD. To this end, at $\alpha=29^{\circ}$, we first calculate $\Delta F^{\ddagger}$ in the temperature range from 300 K to 375 K by $\Delta F^{\ddagger}=F(w_{0}^{\ddagger})+k_{\mathrm{B}}T\ln{({\Delta w_{0}}^{-1}\int_{\mathrm{R}}\mathrm{d}w_{0}\,\mathrm{e}^{-\beta\Delta F(w_{0})})}$, where $\Delta w_{0}$ is the difference between the transition state value $w_{0}^{\ddagger}$ and the lower bound of $w_{0}$ sampled from the equilibrium ensemble of the inverted state. The resultant $\Delta F^{\ddagger}$ varies from 48.196 meV to 49.775 meV, showing that it is approximately constant over the temperature range considered.
@@
Figure 5: (a) For the GNR system at $\alpha=29^{\circ}$, $\ln{(k\,T^{-\frac{1}{2}})}$ is plotted versus $1/T$ to reveal the temperature dependence of the transition rate constant $k$ computed from unbiased simulations. This plot is reminiscent of the Eyring plot [5], but features a different prefactor temperature dependence. Based on Eq. (4), $\Delta F^{\ddagger}$ can be inferred from the linear fit slope. (b) Mean escape time $\tau_{e}$ for the forward transition as a function of temperature, estimated using Eq. (3) for the system at $\alpha=24^{\circ}$.

### v3

# Thermally Activated Snap-through Transitions Controlled by Tunable Metastability

@@ Abstract
The effects of thermal fluctuations on the morphology of two-dimensional materials are hard to harness. We propose that a geometrically constrained graphene nanoribbon (GNR) can exhibit thermally activated snap-through transitions with a predictable and controllable transition rate constant. The energetics and kinetics of such transitions can be fully captured by combining enhanced sampling methods and generalized transition state theory. Using well-tempered metadynamics, we determine the free energy landscape and a pair of asymmetric transition pathways of the GNR system. Notably, generalized transition state theory accurately captures how the transition rate constant responds to temperature and the tunable free energy landscape of our system. This work offers a theoretical framework for elastic metastability, introduces rare event methods into thermalized nanomechanical systems, and provides potential applications in designing nanoscale thermal switches and thermal actuators.
@@
Thermally activated transitions underlie a great variety of processes in physics, chemistry, and biology, with examples ranging from magnetization reversal [myers2002thermally, brown2001langevin] and chemical reactions [hanggi1990reaction, peters2017reaction] to crystal nucleation [wolde1997enhancement] and protein folding [s1994does]. A common characteristic of these systems is the presence of a free energy landscape with metastable states, between which transitions are driven by thermal fluctuations. We propose that elastic structures associated with snap-through transitions [gomez2017critical, radisson2023elastic, wang2024transient], when miniaturized to the nanoscale, constitute an underexplored class of systems where metastable free energy landscapes can emerge and give rise to thermally activated morphological transitions.
@@
The morphological snap-through transition occurs when a bistable elastic structure rapidly switches from one equilibrium state to the other. These transitions allow the generation of fast movements through the release of stored elastic energy and are manifested in various biological and engineered systems, including the Venus flytrap [forterre2005venus], hummingbird beaks [smith2011elastic], mechanical metamaterials [silverberg2014using], and nanoelectromechanical switches [loh2012nanoelectromechanical]. In macroscopic systems, a snap-through transition is typically realized either by applying an external load that overcomes the energy barrier, or by imposing boundary actuation that destabilizes one of the equilibrium states [gomez2018ghosts]. However, due to the presence of thermal fluctuations in microscopic systems, such external interventions may not be required for a snap-through transition to occur. If a nanoscale elastic structure exhibits a free energy barrier that can be surmounted by thermal fluctuations on physically relevant timescales, snap-through transitions would occur spontaneously with an associated rate constant, rendering the system metastable. In this work, we design a hydrogen-terminated graphene nanoribbon (GNR) system subject to adjustable geometric constraints that mimic macroscopic snap-through configurations. This nanoscale setup gives rise to a complex free energy landscape with tunable metastability, thereby enabling thermally activated snap-through transitions to occur on controllable timescales. The resulting system can serve as a prototypical platform for exploring rare event dynamics in thermalized elastic nanostructures.
While early studies reported various effects of thermal fluctuations on the morphology of two-dimensional (2D) materials such as freestanding graphene [fasolino2007intrinsic, los2009scaling, neek2014thermal, havsik2018quantum], only recently have spontaneous transitions with signatures of thermal activation attracted increasing attention [granato2023dynamics, hanakata2024vibrations, granato2025thermal]. The behavior of such microscopic systems coupled to a thermal environment often entails a level of complexity that merits more nuanced scrutiny [jarzynski2017stochastic]. In particular, an accurate description of energetics and kinetics for activated morphological transitions typically relies on capturing the detailed free energy landscape and transition pathways [peters2017reaction, bussi2020using]. For the geometrically constrained GNR system under consideration, enhanced sampling methods, such as metadynamics [barducci2008well], are essential for resolving the GNR’s complex free energy landscape. We propose that rare event approaches, specifically the combination of enhanced sampling techniques and generalized transition state theory [schenter2003generalized], are well suited for the theoretical and computational handling of thermally activated transitions in microscopic elastic systems, even without any prior assumption about system energetics.
Enhanced sampling methods, which allow free energy calculations [frenkel2023understanding, bussi2020using] and accelerated molecular dynamics of rare events [voter1997method, voter1997hyperdynamics, tiwary2013metadynamics, ray2023kinetics], have found widespread applications in chemical, biological, and material systems. Understanding the behavior of metastable nanomechanical systems coupled to thermal environments poses computational challenges analogous to those encountered in previous applications. One would thus expect enhanced sampling to be a natural tool for investigating metastable nanomechanical systems, yet its use in such scenarios remains an open frontier. Here we show that enhanced sampling methods can be applied to the metastable nanomechanical system of interest to derive a well-defined Landau free energy profile, which in turn enables the use of generalized transition state theory to describe and predict the transition kinetics. The accuracy and robustness of this combined approach are demonstrated through analysis of the GNR’s behavior under a tunable free energy landscape. This methodology, long established in statistical mechanics and physical chemistry, can thus be extended to emerging activated processes in microscopic elastic systems, facilitating the design of nanoscale thermal switches and actuators.
Results
Thermally Activated Snap-through Transitions
@@ Results
A GNR can serve as a nanoscale counterpart of elastic strips, which have been increasingly adopted as a canonical system for analyzing the dynamics and mechanisms of snap-through transitions in bistable structures [gomez2017critical, gomez2018pull, pandey2014dynamics, sano2017slip, sano2018snap, radisson2023elastic, radisson2023dynamic, wang2024transient]. Here we consider a zigzag GNR with a length of 15.23 nm and N = 16 zigzag chains, whose edges are hydrogen-terminated to eliminate intrinsic mechanical instabilities [bets2009spontaneous, nikiforov2014formation]. Given the large length and time scales involved in the processes of interest, molecular dynamics (MD) simulations with empirical interatomic potentials constitute the only viable and effective computational approach that retains atomistic resolution [penev2021theoretical, dumitrica2010trends]. The second generation reactive empirical bond order (REBO-II) potential [brenner2002second] is adopted to describe our hydrogen-terminated GNR system.
@@
The GNR can be transformed into a bistable structure when subject to geometric constraint at the boundaries. To achieve this, the outermost seven rows of atoms at each end are rigidly held, leaving a flexible segment of length $L$ = 13.75 nm. A fixed end-shortening $\Delta L=0.08L$ is applied between the two ends, which are then rotated in opposite directions by a small adjustable angle $\alpha$, causing the GNR to adopt one of two mechanical stable states: a natural state and an inverted state, as shown in Fig. 1(a). The equilibrium configurations of these states can be analyzed with Euler–Bernoulli beam theory [gomez2017critical, radisson2023elastic, wang2024transient] in the limit of small transverse displacement $w(x,t)$,
@@
where $\rho$, $b$, and $h$ are the effective density, width, and thickness of the GNR, respectively; $B$ denotes the bending stiffness, and $F$ the applied compressive force. The geometric constraint enters as the boundary conditions and the end-shortening confinement condition [gomez2017critical]. The time-independent solutions of Eq. (1) capture the equilibrium behavior of the system, which can be summarized in a bifurcation diagram in Fig. 1(b). The midpoint displacement $w_{0}$ of the GNR, weighted by $L$, is plotted as a function of the dimensionless control parameter $\mu=\alpha\sqrt{L/\Delta L}$. As $\mu$ increases, the inverted state loses its stability at $\mu^{*}=2$, where it intersects an asymmetric unstable equilibrium branch (denoted as $\mathcal{A}$ state) in a subcritical pitchfork bifurcation. In a macroscopic system, $\mu^{*}$ corresponds to the onset of snap-through towards the natural state. As the inverted state further approaches $\mu^{*}_{2}\approx 2.012$, it undergoes a saddle-node bifurcation and vanishes. Meanwhile, the natural state persists stably over the full range of $\mu$. Note that the equilibrium configurations are entirely controlled by the geometric parameter $\mu$ and are independent of the material parameters of the system [gomez2018ghosts].
Figure 1: (a) Representative configurations of the GNR corresponding to the natural state (top), the asymmetric unstable equilibrium ($\mathcal{A}$ state, middle), and the inverted state (bottom). The rigidly constrained atoms at the two ends impose a fixed end-shortening $\Delta L$ and a symmetric tangent angle $\alpha$ at the boundaries. (b) The bifurcation diagram for the GNR system under consideration. Based on the equilibrium solutions of the Euler–beam equation, the midpoint displacement $w_{0}$ is plotted as functions of the geometric control parameter $\mu$, with solid lines indicating the natural state (violet upper branch) and the inverted state (blue lower branch), and the orange dashed line indicating the $\mathcal{A}$ state. Diamond markers represent the equilibrium averages of $w_{0}$ computed from MD simulations.
To compare the behavior of the GNR in MD simulations with the description of the Euler-beam equation, we monitor $w_{0}$ of the GNR at different values of $\mu$, by setting $\alpha$ to successive integer degrees. For each $\alpha$, the GNR is fully thermalized in the natural and the inverted states separately, and a 5 ns trajectory is initiated from each state to calculate the time average of $w_{0}$. A Langevin thermostat is used to provide the thermalizing environment at 300 K. The time-averaged $w_{0}$ from our MD simulations is found to be in good agreement with the Euler–beam equation, as shown in Fig. 1(b). Naturally, the GNR configuration does not remain stationary at the mechanical equilibrium position of the natural or inverted state, but instead fluctuates around it, thereby defining a thermodynamic equilibrium state. When $\alpha$ exceeds $30^{\circ}$ ($\mu\approx 1.851$), the GNR cannot persist in the vicinity of the inverted state throughout a 5 ns trajectory: it snaps to the natural state under the influences of thermal fluctuations that are not captured by Eq. (1). This indicates that the thermal stability of the inverted state has already been compromised, even though the system is well below $\mu^{*}$, the threshold of mechanical instability.
@@
where $\delta w_{0}(t)=w_{0}(t)-\langle w_{0}\rangle$, and $\langle\cdot\rangle$ denotes the ensemble average in the inverted state. The profile of $\widetilde{C}(t)$ is presented in Fig. 2(b). The time integral of $\widetilde{C}(t)$ yields the dominant local relaxation time scale $\tau_{s}\approx$ 0.32 ns. In addition, we analyze the distribution of escape times in the 100 trajectories. The mean escape time is estimated using maximum likelihood estimation (MLE), yielding a value of $\tau_{e}=51.70$ ns, which is two orders of magnitude greater than the local relaxation time $\tau_{s}$. The cumulative distribution function (CDF) of the escape times is shown in Fig. 2(c), along with the fitted exponential distribution obtained from MLE. A Kolmogorov–Smirnov test produces a $p$-value of 0.957, confirming that the escape times follow an exponential distribution. The separation of time scales ($\tau_{e}\gg\tau_{s}$) [hanggi1990reaction, peters2017reaction, chandler1978statistical] and the exponential distribution of escape times each demonstrate that the snap-through transitions in the GNR constitute a well-defined Poisson process.
Free Energy Landscape and Transition Pathways
@@
As shown in Fig. 2(a), even while remaining within the local equilibrium states, the system exhibits substantial fluctuations in the atomistic potential energy, which obscures a clear description of the transition energetics. Moreover, the system behavior is subject to entropic effects, such as the configurational entropy [palazzesi2017conformational]. Therefore, instead of relying on elastic energy, a free-energy-based approach becomes necessary to capture the thermally activated snap-through at the nanoscale. Here, to determine the free energy profile of the system, we utilize well-tempered metadynamics (WT-MetaD) [barducci2008well], an enhanced sampling method that has been extensively used in complex molecular systems to accelerate rare events and accurately calculate free energy [bussi2020using, dama2014well].
@@
2D WT-MetaD simulations are performed by biasing two carefully chosen collective variables: Since $w_{0}$ characterizes the progression of the snap-through transitions, it serves as the natural order parameter; in addition, motivated by the important role of symmetry breaking in macroscopic elastic strips [radisson2023elastic, wang2024transient], we propose a measure of asymmetry that is readily amenable to biased sampling in WT-MetaD, defined as the difference in transverse displacement between the quarter-length positions measured from the two ends of the GNR: $\Delta w_{\text{qrt}}=w(L/4)-w(-L/4)$. The free energy surface (FES) for the activated snap-through transitions is calculated in the state space of ($w_{0}$, $\Delta w_{\text{qrt}}$), as shown in Fig. 3(a). Two basins corresponding to the inverted state and the natural state are identified, with local minima aligned with $\Delta w_{\text{qrt}}=0$, indicative of their symmetric nature. The transition pathways between the two states are characterized by the minimum free energy paths (MFEPs) connecting the basins, along which the system transitions between the two symmetric states via highly asymmetric intermediate states. To compare with elastic continuum theory, the stable mechanical equilibrium positions from the Euler-beam equation are plotted on the FES, coinciding with the metastable minima. Furthermore, the asymmetric unstable equilibria ($\mathcal{A}$ state) coincide with a pair of first-order saddle points, indicating their role as the transition state in thermal activations.
Figure 3: (a) Free energy surface (FES) computed from 2D WT-MetaD in the ($w_{0}$, $\Delta w_{\text{qrt}}$) space, and (b) unreweighted histogram from a representative 1D WT-MetaD simulation biasing only $w_{0}$, as projected into the ($w_{0}$, $\Delta w_{\text{qrt}}$) space. The collective variable $\Delta w_{\text{qrt}}$ quantifies the degree of asymmetry of the GNR. The locations of the natural state (red circle), inverted state (blue diamond), and the $\mathcal{A}$ state (yellow triangle) are computed from the Euler-beam equation. The minimum free energy paths (MFEPs) on the FES in (a) are consistent with the transition pathways revealed by 1D WT-MetaD in (b). The results shown correspond to $T$ = 300 K and $\alpha=16^{\circ}$ ($\mu\approx 0.9873$), exhibiting qualitative features common across the computed range $0\leq\mu\lesssim 1.789$.
One can determine the one-dimensional (1D) Landau free energy profile along the order parameter $w_{0}$ by integrating the Boltzmann weight over $\Delta w_{\text{qrt}}$ [peters2008path], or equivalently, by WT-MetaD simulations biasing only $w_{0}$, as long as the 1D WT-MetaD trajectories faithfully track the MFEPs on the 2D free energy landscape (as confirmed by Fig. 3). Either way, it is essential to capture the pair of asymmetric pathways in order to correctly characterize the transition energetics.
Corresponding to the transitions recorded in the unbiased simulations at $\alpha=29^{\circ}$ under $T$ = 300 K, the free energy profile $F(w_{0})$ calculated from 1D WT-MetaD is presented in Fig. 4(a), exhibiting two local minima corresponding to the inverted state and the natural state. The forward transition from the inverted to the natural state is associated with a free energy barrier of $F^{+}$ = 71.823 meV, while the backward transition corresponds to a much higher barrier of $F^{-}$ = 2526.776 meV. The substantial height of $F^{-}$ suggests that the reverse transition is kinetically suppressed, as will be addressed in our later analysis. Since the energetics of the system can be controlled by the geometric parameter $\mu$, we have further computed the free energy barriers from WT-MetaD at various values of $\mu$, as shown in Fig. 4(b). When $\mu=0$, the forward and backward transition barriers are equal, implying no energetic preference between the two states. As $\mu$ increases, the barriers become increasingly unequal, showing that the metastability of the GNR system is tunable.
Generalized Transition State Theory in Terms of the Landau Free Energy
Reaction rate theory has been widely employed to describe the kinetics of activated rate processes [hanggi1990reaction, peters2017reaction], prominently including Kramers’ theory [mel1991kramers] and transition state theory (TST) [truhlar1996current, bao2017variational]. In particular, generalized transition state theory in terms of the potential of mean force [schenter2003generalized] (or Landau free energy, depending on the context [ciccotti2025foundations]) provides a rigorous framework for characterizing rare event transition rates when used in conjunction with free energy calculation methods [watney2006calculation]. One can adopt this combined approach to comprehend the kinetics of thermally activated snap-through transitions.
Figure 4: (a) Free energy profile computed from WT-MetaD (five 500 ns production runs) along the order parameter $w_{0}$ for the GNR constrained at $\alpha=29^{\circ}$ ($\mu\approx 1.789$) under temperature $T$ = 300 K. Inset: Magnified view of the free energy profile highlighting the forward transition barrier $F^{+}$. (b) Forward and backward transition barriers, $F^{+}$ and $F^{-}$, as functions of the dimensionless geometric control parameter $\mu$.
@@
Provided that there exists a collective variable $\xi$ such that the transition state dividing surface $\xi(\mathbf{r})=\xi^{\ddagger}$ separates the reactant region $\{\mathbf{r}|\xi(\mathbf{r})<\xi^{\ddagger}\}$ from the product region $\{\mathbf{r}|\xi(\mathbf{r})>\xi^{\ddagger}\}$ in the configuration space, the transition rate constant can be calculated by the reactive flux through this surface. For a system characterized by the 1D Landau free energy $F(\xi)$, the generalized TST rate constant takes the form
$k_{\mathrm{TST}}$ $=\frac{\langle{Z_{\xi}}^{\frac{1}{2}}\rangle_{\xi=\xi^{\ddagger}}}{(2\pi\beta)^{\frac{1}{2}}}\frac{\mathrm{e}^{-\beta F(\xi^{\ddagger})}}{\int_{\xi<\xi^{\ddagger}}d\xi\,\mathrm{e}^{-\beta F(\xi)}}$ (3)
@@
where $\langle\cdot\rangle_{\xi=\xi^{\ddagger}}$ denotes the ensemble average sampled at the transition state and $\int_{\xi<\xi^{\ddagger}}$ denotes the integral over the reactant region. Here $\beta$ is the inverse temperature, $\Delta\xi$ is given by $\Delta\xi=\int_{\xi<\xi^{\ddagger}}d\xi$, $\Delta F^{\ddagger}$ is the free energy of activation, given by
$\Delta F^{\ddagger}=F(\xi^{\ddagger})+k_{\mathrm{B}}T\ln{({\Delta\xi}^{-1}\int_{\xi<\xi^{\ddagger}}d\xi\,\mathrm{e}^{-\beta F(\xi)})},$ (5)
and $Z_{\xi}$ is the inverse effective mass associated with $\xi$, defined as
@@
with $M_{i}$ being the atomic mass associated with the $i$th component of $\mathbf{r}$. For the GNR system under investigation, the order parameter $w_{0}$ serves the role of $\xi$. It should be noted that the free energy of activation $\Delta F^{\ddagger}$ is not identical to the transition barrier $F^{+}$ due to a statistical distribution of the reactant state [schenter2003generalized].
@@
Figure 5: (a) For the GNR system constrained at $\alpha=29^{\circ}$, $\ln{(k\,T^{-\frac{1}{2}})}$ is plotted versus $1/T$ to reveal the temperature dependence of the transition rate constant $k$ computed directly from unbiased simulations. Based on Eq. (4), $\Delta F^{\ddagger}$ can be inferred from the slope of the linear fit. (b) Transition rate constant $k$ for the forward transition as a function of temperature, calculated using Eq. (3) and WT-MetaD for the system constrained at $\alpha=24^{\circ}$.
We justify the applicability of generalized TST to our system by examining the consistency between the free energy of activation $\Delta F^{\ddagger}$ inferred from Eq. (4) based on the transition rates determined in unbiased simulations at different temperatures and the value of $\Delta F^{\ddagger}$ calculated directly from $F(w_{0})$ given by WT-MetaD. To this end, when the system is constrained at $\alpha=29^{\circ}$, we first calculate $\Delta F^{\ddagger}$ in the temperature range from 300 K to 375 K by Eq. (5). The resultant $\Delta F^{\ddagger}$ varies from 48.817 meV to 49.775 meV, showing that it is approximately constant over the temperature range considered.